OpenStudy (anonymous):

Use the order of operations to simplify the expression. –18 + (–4) + 20 + (–3) would be (-4*-3) = 12 + -18 = -6 = 14?

OpenStudy (anonymous):

What I got was (-4*-3) = 12 + -18 = -6 = 14

OpenStudy (anonymous):

there is no multiplication in –18 + (–4) + 20 + (–3)

OpenStudy (anonymous):

I thought if you took the (-4) and the (-3) doing the Order of operations you would multiply them?

OpenStudy (anonymous):

it is 3 additions you do them one at a time from left to right –18 + (–4) + 20 + (–3) = \(-22 +20+(-3)=-2+(-3)=-5\)

OpenStudy (anonymous):

since you have plus signs between all the numbers, it is addition \((-5)+(-4)\) is addition, whereas \((-5)(-4)\) is multiplication

OpenStudy (anonymous):

That makes a lot more sense then. Thanks! I actually understand it now.